# Emergence and Localisation of Semantic Role Circuits in LLMs

### v1

@@ 3.3 Cross-Model Similarity
We assess whether circuits transfer across models by comparing circuit structure and function via graph overlap and spectral geometry.
@@
$J(V_{i},V_{j})=\frac{|V_{i}\cap V_{j}|}{|V_{i}\cup V_{j}|},\quad J(E_{i},E_{j})=\frac{|E_{i}\cap E_{j}|}{|E_{i}\cup E_{j}|}.$ (10)
All attributions are normalised by total absolute mass before thresholding to ensure comparability across model scales.
@@
To compare higher-order circuit geometry, we compute root-mean-square deviation between the smallest $k=16$ eigenvalues of the edge-weighted Laplacian using top-K=50K{=}50 edges:
$d_{\text{spec}}(i,j)=\sqrt{\frac{1}{k}\sum_{m=1}^{k}(\lambda^{(i)}_{m}-\lambda^{(j)}_{m})^{2}}.$ (11)
Smaller $d_{\text{spec}}$ indicates similar information-flow geometry despite differing edge sets. Full spectral analysis methods and Laplacian construction appear in Appendix C.4.
@@ 4 Experimental Setup
4.1 Models
We select models to satisfy three requirements: (i) dense training checkpoints to trace circuit formation across training, (ii) variation in model scale to assess circuit transferability, and (iii) architectural diversity to distinguish general computational strategies from implementation-specific artefacts. Pythia (14M, 410M, 1B; biderman2023pythia) provides full checkpoints from initialisation to convergence, enabling fine-grained analysis of when semantic-role circuits first emerge and consolidate. To test whether these circuits persist beyond a single family, we include LLaMA-1B (touvron2023llama), which differs in tokenisation, training data, and architectural choices. This pairing separates training-time dynamics (within family) from architectural robustness (across families). Although not intended as an exhaustive model survey, this combination is sufficient to assess whether the mechanisms we identify reflect general training-driven pressures rather than results of a specific implementation.
4.2 Datasets and Experimental Software
We instantiate the role-conditioned continuation task (Sec. 3.1) focusing on four roles: Beneficiary, Instrument, Location, and Time. These roles are selected to span different types of semantic relations commonly represented in frame-semantic and role-labelling resources: participant roles (Beneficiary, Instrument) and circumstantial roles (Location, Time). This selection provides variation in lexical and syntactic realisation while keeping the experimental design tractable. Following Sec.3.1, the generated pairs follow token parity and were filtered to retain only examples where the model correctly predicts the role-appropriate target in both contexts. After filtering with Pythia-1B, we obtain 4740 pairs ($\geq$350 per role). Additional roles (Goal, Path, Source, Topic) are analysed in Appendix F.3, where we show that they display the same qualitative localisation and emergence patterns, supporting the generality of our conclusions. Full construction details, filtering procedures, and dataset statistics are provided in Appendix B, and we also provide software and computation specifications in Appendix D.1.
@@ 5 Results
We address three research questions about semantic-role circuits across four roles (Beneficiary, Instrument, Location, Time) in Pythia (14M, 410M, 1B) and LLaMA-1B. Unless otherwise noted, we report results for Pythia-1B. Full results and metric definitions appear in Appendices F and E.
@@ 5.1 Localisation of Role Circuits (RQ1)
We find highly localised circuits for all tested roles, with the top 20 nodes capturing 89–94% of attribution mass at convergence (Table 1). Despite this shared sparsity, roles exhibit qualitatively distinct architectures and developmental trajectories.
@@
Causal-flow analysis (Fig. 2) reveals systematic cross-role differences reflecting semantic demands. For example, Beneficiary exhibits the most complex developmental trajectory: at step 32 (23 nodes), it recruits moderate early-layer attention (a0.h1, a0.h2, a0.h4–a0.h6) alongside extensive MLP branching. By step 71000 (21 nodes), the circuit reorganises into complex late stage architecture with three distinct value-composition operations (a9.h2+V, a14.h5+V, a15.h5+V), indicating intensive feature extraction. At convergence (step 143000, 22 nodes), it maintains rich connectivity with six attention heads. In contrast, other roles converge to simpler architectures: Time achieves mostly MLP implementation by mid-training; Instrument stabilises to balanced hybrid (20 nodes, 4 heads); Location contracts from mid-training expansion to distributed integration. Further, cross-role comparison of top-$K$ components (Fig. 4) shows consistently low Jaccard overlap, confirming that these architectural patterns correspond to distinct, role-specific subgraphs rather than shared motifs. Full circuit evolution for all eight roles appears in Appendix F.4.
⟦t⟧ Role (final step = 143k) Top–20 mass Gini (mass) Beneficiary 0.906 0.439 Instrument 0.917 0.489 Location 0.897 0.519 Time 0.935 0.468
Table 1: Sparsity statistics for role circuits at convergence (Pythia–1B, step 143k). The table reports Top–20 mass (the fraction of attribution carried by the 20 highest-mass nodes) and the corresponding Gini coefficients. All roles exhibit high concentration, indicating that semantic-role computation is mediated by compact subgraphs. Gini values (higher = tighter concentration) show that Time has the most compact attribution profile, while Location spreads mass more evenly across its top components, reflecting small but systematic differences in circuit concentration across roles.
⟦h!⟧ Refer to captionwidth=⟧figures/causal_flow_beneficiary.png
Figure 2: Evolution of the Beneficiary circuit across training. The circuit undergoes complex reorganisation from early exploration (step 32) through intensive mid-training feature extraction (step 71000) to its final architecture (step 143000). Beneficiary’s delayed structural consolidation ($t_{\mathrm{cons}}=2{,}000$) and persistent complexity reflect the computational demands of distinguishing benefactive from alternative role readings. Edge colour encodes operation type (blue: residual flow; green: value-composition; dashed: negative); edge width is proportional to attribution magnitude.
Quantitative localisation.
Table 1 summarises sparsity at convergence. All roles show tight concentration, with the top 20 nodes capturing 89–94% of total mass. Time is the most compact (93.5%), while Location distributes mass slightly more evenly across its top components (Gini = 0.519), reflecting modest differences in how sharply each role concentrates attribution within its circuit.
⟦ht⟧ Refer to captionwidth=0.95⟧figures/role_evolution_dashboard_3.png
Figure 3: Structural and functional dynamics of role circuits across training. Faithfulness (left) shows pronounced role-dependent volatility, with early peaks and mid-training drops, indicating that circuit usefulness does not increase monotonically. In contrast, structural metrics evolve more smoothly: circuit size (middle) contracts gradually, and edge density (right) rises or stabilises over time. These trends show that circuit structure consolidates early and steadily, while functional engagement remains variable across roles.
@@
Temporal analysis (Fig. 3) shows heterogeneous emergence dynamics. Beneficiary displays the smoothest progression: faithfulness increases steadily, and edge density rises steadily (0.09 $\rightarrow$ 0.14), indicating progressive consolidation into compact circuits. Instrument starts with high faithfulness (peak $\simeq$2.45 at step 0), then undergoes a sharp decline and stabilises at lower levels, with edge density dipping before rising, suggesting late consolidation into denser pathways. Location exhibits moderate initial faithfulness with mid-training fluctuations and mild late recovery, while edge density grows steadily, consistent with gradual strengthening of connectivity. Time shows the strongest non-monotonicity: faithfulness begins high ($\simeq$2.0), drops sharply, and partially recovers. Circuit size decreases early but re-expands in the final 70K steps, and edge density peaks highest of any role ($\simeq$0.20 around step 70k) before settling, yielding a distinctive “expand–contract” pattern.
width=⟧figures/role_overlap_heatmaps.png
Figure 4: Cross-role overlap of high-importance components over training (Pythia–1B). Each heatmap shows the Jaccard similarity between the nodes for different roles at a three sample training stages. Overlap remains consistently low, indicating that roles recruit largely distinct component sets. This specialisation strengthens over training, supporting the view that role circuits differentiate rather than collapse into a shared mechanism.
@@
(i) Highly concentrated circuits with role-specific structure. All roles converge to compact circuits (89–94% mass in 20 nodes; Table 1) but exhibit distinct patterns: Instrument uses bridge-augmented MLP spines, Time maintains mid-layer cores, Location densifies gradually, and Beneficiary streamlines progressively. Low cross-role overlap (Fig. 4) confirms strong component specialisation. (ii) Continuous refinement, not discrete transitions. Structural metrics (Top-$K$ mass, edge density, circuit size; Fig. 3) evolve smoothly throughout training. Even roles with large functional swings (Instrument, Time) show gradual consolidation, indicating incremental optimisation rather than abrupt phase shifts. (iii) Structure–function dissociation. Faithfulness is often non-monotonic (e.g., Time drops then recovers) even as sparsity increases. Circuits can be structurally well-formed yet temporarily underutilised, challenging assumptions that presence implies consistent engagement. (iv) Role-specific stabilisation timelines. Roles consolidate at different rates: Instrument and Location stabilise earliest (high Top-$K$ Jaccard by 10K steps), Time stabilises later and less strongly, and Beneficiary consolidates slowly but consistently.
@@ 5.2 Emergence Dynamics (RQ2)
We find that role-binding circuits emerge via continuous refinement rather than discrete phase transitions. Roles become causally indispensable early in training (within 0–128 steps), but structural consolidation unfolds over tens of thousands of steps, producing pronounced structure–function dissociation.
Early indispensability with role-specific emergence.
Roles become causally indispensable at heterogeneous stages (Table 2). Time circuits are indispensable from initialisation ($t_{\text{ind}}=0$), Instrument by step 32, and Location by step 128. Beneficiary presents a distinct pattern: the circuit faithfulness improves throughout training, but ablation never exceeds our conservative $\mu+\sigma$ threshold, leaving $t_{\text{ind}}$ undefined; the circuit is helpful but not provably necessary under our criterion. This structure–function dissociation (rapid structural consolidation with variable functional engagement for most roles; delayed consolidation without strict necessity for Beneficiary) likely reflects differences in cue salience, distributional frequency, and the computational complexity required for disambiguation. We study circuits heterogeneity and temporal instability in Appendix F.4, which provides detailed architectural analyses among roles circuits.
⟦t⟧ Role $t_{\mathrm{ind}}$ (steps) $t_{\mathrm{cons}}$ (steps) Beneficiary 2,000 Instrument 32 8 Location 128 8 Time 0 8
Table 2: Emergence timings for role circuits. For each role, $t_{\mathrm{ind}}$ marks the earliest step where ablation consistently harms performance, and $t_{\mathrm{cons}}$ the first point where the circuit’s structure stabilises (Jaccard $\geq 0.6$ over two steps). All roles consolidate within $\sim$2k steps, but indispensability varies widely, with some circuits becoming necessary immediately and others only much later or not at all.
@@
Piecewise linear regression on Top-$K$ mass trajectories yields change-point estimates with bootstrap confidence intervals spanning multiple orders of magnitude (Table 3). Beneficiary, Instrument, and Location share $\hat{t}_{c}=512$ with CIs ⟦32, 10,000⟧, ⟦8, 71,000⟧, and ⟦8, 7,000⟧, respectively; Time shows $\hat{t}_{c}=7{,}000$ with CI ⟦32, 50,000⟧. These intervals are far too wide to localise any discrete transition: structural sparsification evolves continuously, contrasting with grokking (power2021grokking; nanda2023progress), where change-points are precise and abrupt. Structural metrics evolve smoothly (Fig. 3): Top-$K$ mass reaches $\approx 0.90$ by step 1k and then saturates gradually; edge density increases steadily, with Beneficiary and Instrument showing monotonic growth and Location/Time exhibiting mild “expand–contract” patterns as weak connections are pruned. Final concentration is extreme: Top-20 nodes capture 89.7–93.5% of mass (Table 4), with 95% coverage in just 22–28 nodes.
⟦t⟧ Role $\hat{t}_{c}$ (steps) 95% CI Beneficiary 512 ⟦32, 10,000⟧ Instrument 512 ⟦8, 71,000⟧ Location 512 ⟦8, 7,000⟧ Time 7,000 ⟦32, 50,000⟧
Table 3: Change-point estimates for Top-$K$ node mass obtained via two-segment piecewise linear regression with bootstrap resampling. The extremely wide confidence intervals indicate that structural sparsification evolves gradually and continuously rather than through a discrete phase transition; no statistically localisable break-points appear in the trajectories.
@@
Although structural metrics evolve smoothly, faithfulness trajectories exhibit pronounced non-monotonicity (Fig. 3, left). Beneficiary crashes early then recovers; Instrument and Location peak early then plateau; Time shows the sharpest mid-training crashes and late recovery. This functional volatility despite structural stability shows that circuits can be structurally present while being temporarily underutilised. Structural consolidation precedes functional indispensability by many of steps (Table 2: $t_{\text{cons}}=8$ vs. $t_{\text{ind}}=32$–128), demonstrating that circuit presence does not guarantee circuit engagement.
⟦t⟧ Role Top-5 Top-10 Top-20 $k$ for 80/90/95% Beneficiary 0.395 0.611 0.906 16 / 20 / 25 Instrument 0.492 0.688 0.917 14 / 19 / 23 Location 0.483 0.673 0.897 15 / 21 / 25 Time 0.386 0.635 0.935 15 / 18 / 22
Table 4: Final-step concentration of role circuits. Top–$K$ mass and coverage thresholds reveal that each role exhibits a sharply concentrated attribution distribution. The small $k$ values required to cover most of mass show that a limited subset of nodes drives the majority of the circuit’s computation, with role-to-role variation reflecting consistent but distinct concentration patterns.
@@
(i) Continuous refinement throughout training. Structural metrics (Top-$K$ mass, edge density, circuit size) evolve smoothly with no discrete transitions; wide change-point CIs indicate no identifiable phase shifts. (ii) Early indispensability, prolonged consolidation. Roles become indispensable within 0–128 steps, but structural refinement continues for 140K+ steps. Structure stabilises early, but attribution mass continues to redistribute. (iii) Structure–function dissociation and delayed engagement. Circuits stabilise structurally ($t_{\text{cons}}=8$) many steps before becoming functionally indispensable ($t_{\text{ind}}=32$–128). Faithfulness shows crashes and recoveries even as sparsity increases, illustrating that circuit presence does not mean circuit engagement. (iv) Role-specific developmental trajectories. Despite shared qualitative patterns (continuous sparsification, early consolidation), roles show distinct functional dynamics: Beneficiary recovers after an early crash, Location plateaus mid-training, Instrument peaks then declines, and Time exhibits severe mid-training instability.
⟦ht⟧ Refer to captionwidth=0.8⟧figures/cross_scale_overlap_location.png Refer to captionwidth=0.8⟧figures/cross_scale_overlap_instrument.png
Figure 5: Cross-family correspondence for Location (top) and Instrument (bottom). Average node-(left) and edge-level (right) overlaps between Pythia–1B and LLaMA–1B. Node sets align substantially more than edges, suggesting shared component selection but model-specific routing.

### v2

@@ 3.3 Cross-Model Similarity
We assess circuit transferability across models by comparing structure and function via graph overlap and spectral geometry.
@@
To compare higher-order circuit geometry, we compute root-mean-square deviation between the smallest k=16k{=}16 eigenvalues of the edge-weighted Laplacian using top-K=50K{=}50 edges: $d_{\text{spec}}(i,j)=\sqrt{\frac{1}{k}\sum_{m=1}^{k}(\lambda^{(i)}_{m}-\lambda^{(j)}_{m})^{2}}.$
Smaller $d_{\text{spec}}$ indicates similar information-flow geometry despite differing edge sets (full details in App. D.4).
@@ 4 Experimental Setup
Models.
We select models to satisfy three criteria: (i) dense training checkpoints to trace circuit formation, (ii) variation in scale to assess transferability, and (iii) architectural diversity to distinguish general strategies from implementation-specific artefacts. Pythia (14M, 410M, 1B) (biderman2023pythia) provides comprehensive checkpoints, enabling fine-grained analysis of circuit emergence. To test whether these circuits persist beyond a single family, we include LLaMA-1B (touvron2023llama), which differs in tokenisation, training data, and architectural choices. This pairing separates training-time dynamics (within family) from architectural robustness (across families).
Datasets and Experimental Software.
We instantiate the role-conditioned continuation task (Sec. 3.1) across eight semantic roles spanning participant (Beneficiary, Instrument), directional (Goal, Source, Path), locative and temporal (Location, Time), and propositional (Topic) categories, providing broad coverage of predicate–argument structure types (fillmore1976frame; palmer2005proposition; carnie2021syntax). After filtering to retain only examples where models correctly predict role-appropriate targets in both contexts, we obtain approximately 6,000–8,000 pairs per model across roles (exact counts appear in Table 5). Full dataset construction details, role categorisations, filtering procedures, and statistics are provided in App. B. Software and computation specifications appear in App. C.2.
@@ 5 Results
We address three research questions on semantic-role circuits across eight roles using Pythia (14M, 410M, 1B) and LLaMA-1B. For presentation clarity, we report main text results for four representative roles, Beneficiary, Instrument, Location, and Time; spanning participant, locative, and temporal categories. Results for the remaining roles (Goal, Source, Path, Topic) follow similar qualitative trends and are provided in App. F.3. Unless otherwise specified, results refer to Pythia-1B. Comparisons across model scales and architectures are explicitly noted. Full results and metric definitions are given in App. F and E.
@@ 5.1 Localisation of Role Circuits (RQ1)
We find highly localised circuits for all tested roles, with the top 20 nodes capturing 89–92% of attribution mass at convergence (Tab. 1). Despite this shared sparsity, roles exhibit qualitatively distinct architectures and developmental trajectories.
@@
Causal-flow analysis (Fig. 2) reveals systematic cross-role differences. Beneficiary develops the most complex architecture, combining early attention, extensive MLP branching, and rich late-layer connectivity with multiple value-composition operations. In contrast, Time shifts from early-layer attention to convergent mid-to-late processing; Instrument stabilises to a balanced hybrid circuit; and Location transitions from mid-training expansion to more distributed integration. Low to mid Jaccard overlap across roles (Fig. 4) confirms these correspond to distinct, role-specific subgraphs. Full circuit evolution for additional roles in App. F.4.
Figure 2: Evolution of the Beneficiary circuit across training. The circuit reorganises from early exploration (step 32), through mid-training feature extraction (step 71K), to its final architecture (step 143K). Edge colour encodes operation type ( blue: residual; green: value composition; dashed: negative), and edge width indicates attribution magnitude.
Role T-5 T-10 T-20 $k$ for 80/90/95% Gini (mass)
Beneficiary 0.395 0.611 0.906 16 / 20 / 25 0.439
Instrument 0.492 0.688 0.917 14 / 19 / 23 0.489
Location 0.483 0.673 0.897 15 / 21 / 25 0.519
Time 0.466 0.664 0.910 15 / 20 / 24 0.475
Table 1: Final-step concentration profiles of role circuits (step 143K). For each role, we report cumulative attribution mass captured by the top (T-$k$) nodes (T-5, T-10, T-20), the smallest $k$ achieving 80/90/95% total mass, and the Gini coefficient over attribution mass. Together, these metrics show that role circuits are highly sparse, with a small set of nodes dominating computation, while differences in Gini reflect variation in how evenly mass is distributed beyond the top contributors.
Figure 3: Structural and functional dynamics of role circuits across training. Faithfulness (left) shows pronounced role-dependent volatility. In contrast, structural metrics evolve smoothly: circuit size (middle) contracts gradually, and edge density (right) rises or stabilises, showing structure consolidates early and steadily while functional engagement remains variable.
@@
Temporal analysis (Fig. 3) shows heterogeneous emergence dynamics. Instrument starts with high faithfulness, undergoes a sharp decline, then stabilises with edge density dipping before rising, suggesting late consolidation into denser pathways. Beneficiary exhibits the most pronounced volatility: faithfulness spikes sharply at step 50k, drops sharply, partially recovers; circuit size decreases early but re-expands in the final 70K steps, yielding a distinctive “expand–contract” pattern. Location and Time maintain relatively stable faithfulness after initial drops, with edge density showing moderate fluctuations, consistent with gradual strengthening of connectivity. Circuit sizes contract from initialisation (29–54 nodes) to convergence (31–36 nodes) across all roles, with Beneficiary showing the largest initial size and most rapid contraction.
Figure 4: Cross-role overlap of high-importance components over training (Pythia–1B). Overlap remains consistently low across training stages, indicating roles recruit largely distinct component sets, supporting circuit differentiation rather than shared mechanisms.
@@
(i) Highly concentrated circuits with role-specific structure. All roles converge to compact circuits (89–92% mass in $\leq$20 nodes) with role-specific architectures confirmed by low cross-role overlap. (ii) Continuous refinement, not discrete transitions. Structural metrics (circuit size, edge density) evolve smoothly throughout training, indicating continuous refinement rather than discrete transitions. (iii) Structure–function dissociation. Faithfulness exhibits pronounced non-monotonicity even as circuits structurally consolidate. This demonstrates that circuits can be structurally well-formed yet functionally but temporarily underutilised. (iv) Role-specific stabilisation timelines. Roles show distinct functional trajectories: Beneficiary exhibits extreme volatility, Instrument sharp early decline then stability, Location and Time moderate stability after initial adjustment.
Role $t_{\mathrm{ind}}$ (steps) $t_{\mathrm{cons}}$ (steps)
Beneficiary 50,000
Instrument 32 128
Location 128 128
Time 512 5,000
Table 2: Emergence timings for role circuits. $t_{\mathrm{ind}}$ marks earliest step where ablation consistently harms performance; $t_{\mathrm{cons}}$ marks when structure stabilises. All roles consolidate within $\sim$2k steps, but indispensability varies widely.
@@ 5.2 Emergence Dynamics (RQ2)
We find that role-binding circuits emerge via continuous refinement rather than discrete phase transitions. Roles become causally indispensable early (0–512 steps), but structural consolidation unfolds over tens of thousands of steps, producing pronounced structure–function dissociation.
Early indispensability with heterogeneous emergence.
Roles become indispensable at different stages (Tab. 2): Instrument by step 32, Location by step 128, and Time by step 512. Beneficiary exhibits fluctuating improvement in faithfulness and never exceeds our conservative indispensability threshold ($\mu+\sigma$), leaving $t_{\text{ind}}$ undefined and indicating usefulness without provable necessity under this criterion. This dissociation likely reflects differences in cue salience, frequency, and computational demands. Architectural heterogeneity across roles is detailed in App. F.4.
@@
Piecewise linear regression on Top-$K$ mass yields extremely wide change-point confidence intervals (CIs; Tab. 3), ruling out discrete transitions. This observation contrasts with grokking, where change-points are sharp, localised and abrupt (power2021grokking; nanda2023progress). Instead, structural metrics mostly evolve smoothly throughout training (Fig. 3), consistent with gradual sparsification, with some per role specification like Location exhibiting mild “expand–contract” patterns. Final concentration is extreme: Top-20 nodes capture 89.7–91.7% of mass, with 95% coverage in just 23–25 nodes (Tab. 1).
Role $\hat{t}_{c}$ (steps) 95% CI
Beneficiary 10,000 [128, 10,000]
Instrument 50,000 [128, 10,000]
Location 10,000 [128, 10,000]
Time 5,000 [128, 114,000]
Table 3: Change-point estimates for Top-$K$ node mass via piecewise linear regression. Extremely wide confidence intervals indicate gradual, continuous evolution without discrete phase transitions.
@@
Although structural metrics evolve smoothly, faithfulness trajectories exhibit pronounced non-monotonicity (Fig. 3, left). Structural consolidation precedes functional indispensability by several steps (Tab. 2: $t_{\text{cons}}{=}128$ vs. $t_{\text{ind}}{=}32$–512), demonstrating that circuit presence does not guarantee circuit engagement.
Robustness to scaffold variation and frequency effects.
To test whether circuits reflect abstract role binding rather than memorised lexical scaffolds, we evaluate within-role paraphrases for Location and Instrument. Faithfulness and sparsity remain largely unchanged under paraphrase (App. F.5). Moreover, circuit properties at convergence show weak correlations with (i) per-role filtered sample size and (ii) scaffold frequency in the pretraining corpus (App. F.6), suggesting the observed architectural differences are not driven by dataset volume or training-signal strength alone.
@@
(i) Continuous refinement throughout training. Structural metrics evolve smoothly, with wide change-point CIs indicating no discrete phase shifts. (ii) Early indispensability, prolonged consolidation. Roles become indispensable within 0–512 steps, while structural refinement continues for over 143K steps as attribution mass redistributes. (iii) Structure–function dissociation and delayed engagement. Circuits stabilise structurally many steps before becoming functionally indispensable; faithfulness shows crashes and recoveries even as sparsity increases, indicating that circuit presence does not mean engagement. (iv) Role-specific developmental trajectories. Despite shared qualitative patterns, roles exhibit distinct functional dynamics reflecting role-specific computational demands.